OpenStudy (anonymous):

calculate the number of molecules present in 392 gram of sulphuric acid

OpenStudy (abmon98):

\[1 mole--->6.02*10^{23} molecules\] \[4 mole-->x molecules\] \[x=4*6.02*10^{23}/1\] Number of Moles=Mass(g)/Molar Mass Provide each elements atomic weigh S:32 grams H:1 gram O:16 H2SO4 Molar Mass:(2*1)+(32*1)+(16*4)=98 grams Plug in 392 as the mass and 98 as the molar mass 392/98= 4 moles of Sulfuric

OpenStudy (anonymous):

\[molar mass of H2SO4 = 98.086 g/mol\] \[392 g \times \frac{ 1 mole }{ 98.086 g}\times \frac{ 6.02\times 10^{23} }{ 1 mole}= 2.41\times 10^{^{24}} molecules\]